Accelerator Update for June 4, 2008 - June 6, 2008
Wednesday June 4
The day shift began with Operations monitoring Tevatron (TeV) store 6191 (228.7E30) and the Recycler stash (118.0E10), with the Antiproton Source (Pbar) stacking (18.3E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:39 AM, a Linac RF station (LRF2) tripped off. It reset.
At 9:04 AM, the MI Power Supply interlock loop dropped out on a bypass fault. There was no indication of a cause. It reset.
Recycler experts reported at 9:43 AM that they had completed their Pelletron work.
The electrons absorb energy from the antiprotons shrinking the size and spread of the antiprotons and, just like a cooled gas, the antiprotons become denser. This allows the Recycler to store more antiprotons and then, when transferred, increases the luminosity in the Tevatron.
At 12:29 PM, a TeV RF station (TRF7) tripped off.
A TeV RF expert investigated the TRF7 problem at 1:38 PM.
TeV experts reported at 1:41 PM that the quench was due to bad beam positions at TeV sector E1.
A MI expert began a coalescing study at 1:55 PM.
Operations prepared CDF and D0 for access at 1:59 PM.
Operations turned off the Linac beam at 2 PM so that Linac experts could work on a Klystron RF station (KRF6).
Linac experts reported at 2:11 PM that they were looking at two Linac quadrupole power supplies (QPS103 & 104). They were done in 15 minutes.
FESS reported at 2:18 PM that the SY enclosure exhaust fan was up and running.
An RF expert investigated a problem with MIRF3 at 2:48 PM. This only took a few minutes.
A MI expert began a no-beam stacktail measurement at 2:49 PM. Stacking resumed about 10 minutes later.
At 2:49 PM, LRF4 tripped off. It reset.
Operations reported at 4:02 PM that the Recycler had stashed antiprotons four times during the day shift.
CDF came out of access at 4:31 PM.
The cryo and vacuum technicians came out of the TeV access at 4:39 PM.
Operations reported at 7:21 PM that the B1 cryo system was ready to go.
Operators reported at 7:26 PM that they had search and secured SY enclosures C, D, and E.
At 8:06 PM, Operations reported that a Tevatron Electron Lens (TEL2) had tripped off on a bad vacuum status and would not reset. Operations contacted an expert.
Operations put the TeV through a dry squeeze at 8:30 PM.
Operations put the TeV through a wet squeeze at 10 PM.
Operations put the TeV through another wet squeeze at 10:39 PM.
At 10:49 PM, Operations reported that the Booster had been making up a lot of water for its LCW system. Operators searched, but found no leaks.
Operations put the TeV into shot setup at 11:03 PM.
A MI expert began a coalescing check at 11:27 PM.
Operations reported at 11:54 PM that the Recycler had stashed four times during the evening shift.
Thursday June 5
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (90.3E10), with Pbar stacking (10.5E10), and with BNB taking beam.
Operations established store 6195 at 12:25 AM with an initial luminosity of 286.2E30.
At 4:32 AM, Operations reported a heat load on the TeV sector E1 cryo system. There was a cryo valve (CV3D) that should have been opening, but wasn’t responding. Operations called in an expert.
At 5:36 AM, Operations reported that many power supplies at TeV sector F27 had tripped off on a LCW flow indication. Operations called in an expert.
At 5:38 AM, the cryo expert reported that he had tried to force CV3D to open, but it wouldn’t move. The valve seemed to be broken. Cryo experts at CHL overpowered the heat leak with more flow. An access would be needed to repair the valve.
At 7:26 AM, Operations reported a LCW leak was found coming from a Pbar Debuncher power supply (D:H717) and that leak caused the F27 problems. Operations called in power supply experts.
Operations reported at 7:57 AM that the Recycler had stashed four times during the midnight shift.
Two Recycler experts began an Ecool study at 7:59 AM.
At 8:30 AM, a Linac quadrupole power supply (L:QPS103) tripped off. It reset.
A power supply expert reported at 8:36 AM that he had replaced a bad LCW hose for D:H717. Pbar resumed stacking at 8:39 AM.
SY experts began sending beam to the SY dump at 9:09 AM.
At 10:19 AM, Operations reported that stacking went bad for a few minutes and then came back to normal on its own. Operations informed Pbar experts.
A vacuum expert reported at 10:24 AM that the vacuum at TeV sector E1 was going bad.
The SY experts completed their tune at 10:41 AM. Operators disabled the SY beam.
At 1:09 PM, a Linac bulk quadrupole power supply (L:Q3BULK) tripped off. It reset.
At 1:42 PM, the Linac Klystron Vernier tripped off on a water flow fault. It reset.
Operations reported at 1:43 PM that they had intended to end store 6195 in 15 minutes, but the store aborted at 1:50 PM. There was no quench.
At 1:57 PM, a Booster vacuum pump tripped off and caused BRF13 & 14 to also trip off. Everything reset.
Operations prepared CDF and D0 for access at 2:38 PM.
Operations prepared the TeV for access at 2:56 PM.
At 3 PM, Operations reported that they had lost communication with two Recycler correction magnet power supplies (R:V519 and R:H520). Operations contacted an expert.
A Recycler expert reported at 3:35 PM that he had removed the reference cables for two power supplies. He said that the CAMAC 453 cards for these channels needed to be replaced after the store goes in, but before the next stash.
Operations reported at 4:02 PM that the Recycler had stashed four times during the day shift.
At 4:05 PM, CDF dropped their interlocks.
D0 came out of access at 4:21 PM.
The TeV came out of access at 4:35 PM.
A cryo system expert reported at 5:10 PM that they were turning off the TeV sector C4 cold compressor to change a cold valve.
CDF came out of access at 5:17 PM. Operators began a search and secure.
At 5:38 PM, the MI RF South Anode Power Supply tripped off. It reset.
At 5:48 PM, BRF18 tripped off for the first of three times in a little over an hour. It would not reset after the last trip. Operators bypassed the station, lowered the Booster beam intensity, and contacted an expert.
A Booster RF expert reported at 7:42 PM that he had repaired BRF18. Operations raised the Booster beam intensity back to normal.
At 9:06 PM, KRF2 tripped off. It reset.
Operations reported at 9:52 PM that the TeV sector C4 cryo system was ready to go.
Operations put the TeV through a dry squeeze at 10:19 PM.
Operations put the TeV through a wet squeeze at 11:25 PM.
At 11:50 PM, the Linac Klystron Vernier tripped off. It reset.
Operations reported at 11:55 PM that the Recycler had stashed once during the evening shift.
Friday June 6
The midnight shift began with Operations monitoring the stash (339.2E10), with Pbar stacking (149.4E10), and with BNB taking beam.
Operations put the TeV through a wet squeeze at 12:09 AM.
The Run Coordinator reported at 12:25 AM that D0 discovered a problem with their Central Platform RMI that took out ½ of their silicon detector.
Operations prepared D0 for an access at 1:24 AM.
At 1:35 AM, the TeV sector C4 wet engine came into alarm. Cryo system experts investigated and said everything looked okay.
D0 came out of access at 2:14 AM. D0 reported that that the Central Platform RMI had a bad blower. It would take an eight-hour access to repair.
Operations put the TeV through a dry squeeze at 2:58 AM.
At 3:14 AM, the Linac Klystron Vernier tripped off. It reset.
Operations put the TeV into shot setup at 3:20 AM.
At 4:19 AM, Operations reported that the B0 BLMs were pulling the permit. Operators rebooted the B0 BLM system and restored a file.
Operations reported at 4:46 AM that the 2.2-second shot setup timeline was hurting stacking. Operators put Pbar in standby.
Operations established store 6200 at 5:16 AM with an initial luminosity of 298.41E30.
Pbar resumed stacking at 5:38 AM.
Operators replaced the bad Recycler 453 card and reconnected the reference cables at 5:41 AM.
At 7:05 AM, BRF3 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed once during the midnight shift.
The Plans for Friday and the Weekend
The Run Coordinator will decide later today about an eight hour D0 access; if a D0 access is allowed there are a number of other jobs that could be completed in that time frame. Other than the possible access, the plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 06/09/2008 email Fermilab|
|Security, Privacy, Legal|